OpenStudy (anonymous):

Find the area of three-fifths of a circle with radius 15 ft.

OpenStudy (tkhunny):

ALL of the area is what?

OpenStudy (anonymous):

what?

OpenStudy (shamil98):

find the area of the circle first.. then multiply it by 3/5 to get your answer.

OpenStudy (shamil98):

area of circle = ?

OpenStudy (shamil98):

area of circle = pi r^2 so r = 15 (15)^2 pi * 3/5 = ?

OpenStudy (anonymous):

424.11

OpenStudy (anonymous):

i accidentally used circumference instead of area